OpenStudy (anonymous):

The reciprocal of a number equals the reciprocal of the difference of 4 and the number. Find the number

OpenStudy (anonymous):

Please include steps if possible. :)

hero (hero):

1/n = 1/(4-n)

OpenStudy (anonymous):

1/x=1/(4-x)

hero (hero):

cross multiply

OpenStudy (anonymous):

x=4-x 2x=4 x=2

hero (hero):

4-n = n 2n = 4 n = 2

OpenStudy (anonymous):

Thanks you. :)